Bandar Industries Berhad of Malaysia manufactures sporting equipment. One of the company’s products, a football helmet for the North American market, requires a special plastic. During the quarter ending June 30, the company manufactured 3,100 helmets, using 2,232 kilograms of plastic. The plastic cost the company $14,731.

According to the standard cost card, each helmet should require 0.67 kilograms of plastic, at a cost of $7.00 per kilogram.

Required:

1. What is the standard quantity of kilograms of plastic (SQ) that is allowed to make 3,100 helmets?

2. What is the standard materials cost allowed (SQ × SP) to make 3,100 helmets?

3. What is the materials spending variance?

4. What is the materials price variance and the materials quantity variance?
Answer #1

Standard Quantity allowed = 3100*0.67

= 2077

Standard materials cost allowed = 2077*7

= 14,539

Materials spending variance = 14,539 - 14,731

= 192 U

Materials Price Variance = (SP-AP) * AQ

= (7 - 14731/2232) * 2232

= 893 F

Materials Quantity Variance = (SQ-AQ) * SP

= (2077-2232)*7

= 1085 U